# Take History as a Mirror in Heterogeneous Federated Learning

### v1

@@ Abstract
Federated Learning (FL) allows several clients to cooperatively train machine learning models without disclosing the raw data. In practice, due to the system and statistical heterogeneity among devices, synchronous FL often encounters the straggler effect. In contrast, asynchronous FL can mitigate this problem, making it suitable for scenarios involving numerous participants. However, Non-IID data and stale models present significant challenges to asynchronous FL, as they would diminish the practicality of the global model and even lead to training failures. In this work, we propose a novel asynchronous FL framework called Federated Historical Learning (FedHist), which effectively addresses the challenges posed by both Non-IID data and gradient staleness. FedHist enhances the stability of local gradients by performing weighted fusion with historical global gradients cached on the server. Relying on hindsight, it assigns aggregation weights to each participant in a multi-dimensional manner during each communication round. To further enhance the efficiency and stability of the training process, we introduce an intelligent $\ell_{2}$-norm amplification scheme, which dynamically regulates the learning progress based on the $\ell_{2}$-norms of the submitted gradients. Extensive experiments demonstrate that FedHist outperforms state-of-the-art methods in terms of convergence performance and test accuracy.
Introduction
As a distributed machine learning paradigm, Federated learning (FL) (Konečnỳ et al. 2016; McMahan et al. 2017) has garnered widespread attention in recent years due to its ability to protect the privacy of end-user data. In a classical FL workflow, each client uses its private data to train a local model and then uploads the gradient to a central server. The central server aggregates the collected local gradients into a global gradient to update the global model, which is then sent back to the clients for subsequent training rounds. The entire process does not leak the raw data. As such, FL effectively safeguards data privacy and provides unique advantages in various domains including healthcare (Brisimi et al. 2018), autonomous driving (Liang et al. 2022), and mobile keyboard prediction (Hard et al. 2018).
@@ Introduction
A well-known FL aggregation algorithm is federated averaging (FedAvg) (McMahan et al. 2017), where the server aggregates updates from a selected group of local clients and computes the global model by applying a weighted averaging strategy. Due to the system and statistical heterogeneity among devices, this approach often encounters the straggler effect (Cipar et al. 2013), which severely hampers the training progress. In response, (Xie, Koyejo, and Gupta 2019) introduced asynchronous FL (AFL), where the server no longer waits for all clients but can promptly update the global model upon receiving a single local gradient. However, AFL may lead to global model bias towards fast clients, and the frequent data transmissions may result in server crashes (Shi, Zhou, and Niu 2020). Some studies have tried to seek a compromise between synchronous and asynchronous approaches (Wu et al. 2020; Ma et al. 2021; Zhou et al. 2022), where K-async FL (Zhou et al. 2022) has demonstrated promising results, particularly in highly heterogeneous environments. In K-async FL, the server initiates aggregation as soon as it receives the first K local gradients, while the remaining clients continue their local training.
@@
Despite the favorable performance of K-async FL, it is still constrained by two challenges, as shown in Figure 1. On the one hand, the presence of non-independent and identically distributed (Non-IID) data poses a significant challenge in training a well-performing global model (Zhao et al. 2018). On the other hand, the existence of stale local gradients often affects the utility of the global gradient, even leading to divergence (Zhang et al. 2023a). Currently, it is common practice to adopt strategies such as variance reduction (Karimireddy et al. 2020; Li, He, and Song 2021), regularization (Li et al. 2020), and momentum (Li et al. 2019; Xu and Huang 2022) to accommodate heterogeneous scenarios. As for the latter, the mainstream approach involves assigning different weights to local gradients based on their staleness (Chen, Sun, and Jin 2019; Damaskinos et al. 2022) to alleviate the weight divergence. While some algorithms can address both aforementioned issues simultaneously, they either rely on a sole weighted criterion (Zhou et al. 2022) leading to a biased allocation of weights, or necessitate specific prior knowledge that is not readily accessible such as the grouping of clients based on varying computational capacities (Li et al. 2022b), resulting in challenges in practical application.
@@
To address these two challenges, we propose FedHist, a novel K-async FL framework that is capable of handling Non-IID data and stale gradients simultaneously. Specifically, we leverage a historical gradient buffer at the server-side to facilitate knowledge sharing across clients, enhancing the adaptability to Non-IID data. And we use the historical updates of clients to evaluate their utilities, as a factor for determining their weights during the near future aggregation stage. Additionally, we investigate factors that influence convergence and propose a strategy for amplifying the $\ell_{2}$-norms of aggregated gradients adaptively to accelerate the convergence of K-async FL. More importantly, FedHist does not require any prior knowledge about the clients, such as computational capabilities, and maintains the same level of privacy as FedAvg. In summary, with the aid of three core components, FedHist effectively addresses the challenges posed by Non-IID data and stale gradients in AFL. Our main contributions can be summarized as follows:
@@
  • We introduce a gradient buffer to achieve history-aware aggregation on the basis of improving the stability of local gradients and propose a $\ell_{2}$-norm amplification strategy. To the best of our knowledge, we are the first to propose using historical gradients to mirror client utility and expedite the future convergence process.
  • Extensive experiments on public datasets demonstrate that FedHist significantly outperforms state-of-the-art methods. Furthermore, the ablation experiments validate the flexibility and scalability of FedHist.
Related Work and Motivation
Non-IID Data
Statistical heterogeneity refers to the variations in data distribution and data characteristics among individual clients, which has a significant impact on model performance (Chen et al. 2020; Li et al. 2022c). (Zhao et al. 2018) investigates the impact of Non-IID data on FL and observes that as data heterogeneity increases, the accuracy of the classical synchronous algorithm FedAvg diminishes. To mitigate such heterogeneity, some classical approaches involve introducing proximal regularization to reduce variance (Li et al. 2020; Karimireddy et al. 2020; Li, He, and Song 2021). Another intuitive approach is to leverage non-sensitive data to improve training efficiency (Zhang et al. 2022; Zhao et al. 2018) or directly distribute models across clients (Fan et al. 2022). Some studies employ the method of client selection (Lai et al. 2021; Nishio and Yonetani 2019; Li et al. 2022a) with the aim of prioritizing clients that possess higher statistical and system utility. However, these methods either yield only marginal performance improvements or require prior knowledge of the statistical characteristics of each client or rely on additional data, making their implementation challenging in many practical scenarios.
Gradient Staleness
Stale gradients often possess more noise, which may impede convergence speed and even lead to training failure, particularly in Non-IID data scenarios (Dai et al. 2018; Zhou et al. 2018; Zhang et al. 2015). Hence, it has become a consensus in the majority of works to establish a negative correlation between the aggregation weights and the degree of staleness (Jiang et al. 2017; Chen, Sun, and Jin 2019; Su and Li 2022). Also, client grouping (Zhang et al. 2023b; Li et al. 2022d; Wang et al. 2021) can mitigate the impact of the straggler effect. Unfortunately, stragglers may also hold valuable information that could contribute to the model convergence (Liu et al. 2021; Zhou et al. 2022). Hence, relying solely on staleness to assign weights to clients can lead to unfair weight distribution outcomes. Several approaches propose boosting factors to bias the weights towards stale gradients, typically calculated based on the extent of deviation between stale gradients and fresh gradients (Abdelmoniem et al. 2023; Su and Li 2022). (Damaskinos et al. 2022) utilizes the Bhattacharyya coefficient to capture the significance of each gradient, but this approach carries the risk of inadvertently exposing the label distribution of individual client data, potentially leading to privacy breaches.
Trade-off
There exists a fundamental contradiction between simultaneously addressing the issues of Non-IID data and gradient staleness (Zhou et al. 2022). This stems from the fact that resolving the former necessitates considering as many gradients as possible, while the latter requires selecting a minority of low-staleness gradients. These two strategies can be summarized as “more is better” and “quality over quantity”. The interaction among them propels us to propose a novel weighting scheme for aggregation, which involves exploring the utility of diverse gradients beyond staleness, to achieve a more comprehensive information utilization.
Figure 2: The overall framework of FedHist.
Methodology
FedHist tackles the challenges posed by Non-IID data and gradient staleness in K-async FL simultaneously through three fundamental components: enhancement of gradient stability, historical-aware aggregation, and intelligent $\ell_{2}$-norm amplification. Figure 2 illustrates the framework of FedHist.
Problem Statement
Suppose there are $N$ clients, denoted as $c_{1},c_{2},\ldots,c_{N}$. Each client $c_{i}$ trains a model using its local dataset $\mathcal{D}_{i}=\{\xi_{i,1},\xi_{i,2},\cdots,\xi_{i,\lvert{\mathcal{D}_{i}}\rvert}\}$, where $\xi_{i,j}$ represents the $j$-th data point $\{x_{j},y_{j}\}$ on client $i$. Our objective is to collaboratively train a global model $w$ using the entire dataset $\mathcal{D}=\bigcup_{i\in N}\mathcal{D}_{i}$ without compromising the privacy of the raw data. For client $c_{i}$, its empirical loss is defined as:
@@
where $w$ is a parameter vector, which is typically synchronized with the server at the beginning of each round, and $f(\cdot)$ is a loss function. The objective is to solve:
@@
where $K$ represents the number of participating clients in each communication round. In round $r+1$, client $c_{i}$ updates the parameter vector $w_{i}^{r}$ through mini-batch stochastic gradient descent (SGD) to minimize the objective function by Eq. (3):
@@
where $w_{i}^{r+1}$ represents the model of client $c_{i}$ in round $r+1$, $w^{r}$ denotes the global model in round $r$, and $\eta_{i}$ represents the learning rate of client $c_{i}$. We use $g(w_{i}^{r})$ to denote the gradient submitted by $i$-th client in round $r$. Upon receiving $K$ gradients, the server updates the global model parameters by Eq. (4) accordingly:
@@
Here, $w^{r+1}$ represents the global model in round $r+1$, $\eta$ denotes the learning rate of the server, $p_{i}^{r}$ denotes the weight of $i$-th client in round $r$, and $\sum_{i=1}^{K}p_{i}=1$, generally.
Enhancement of Gradient Stability (EGS)
FedHist integrates a buffer at the server-side that dynamically maintains the most recent historical gradients. This buffer is further divided into two fixed-length lists, denoted as $\mathcal{G}$ and $\mathcal{L}$, caching the global gradients and local gradients from the most recent $h$ communication rounds, respectively. And gradients from even earlier rounds will be cleared.
The goal of $\mathbf{EGS}$ is to select an appropriate collaboration gradient from $\mathcal{G}$ for each local gradient uploaded from the client, and then execute a weighted aggregation process. Quantitatively measuring the degree of similarity between two gradient vectors $g(w_{i})$ and $g(w_{j})$ can be achieved by computing their cosine similarity, as shown in Eq. (5):
$cos\langle g(w_{i}),g(w_{j})\rangle=\frac{{g(w_{i})}\cdot{g(w_{j})}}{\|{g(w_{i})}\|\,\|{g(w_{j})}\|}.$ (5)
We select the corresponding collaborative gradient for each local gradient based on the principle of minimizing similarity, as it implies a higher potential of sharing knowledge between two gradients and avoids overfitting (Hu et al. 2022). This fine-grained cross-aggregation method mitigates gradient divergence and rectifies bias in local gradients. Thus the objective is to find the optimal collaborative gradient $g(w_{i,co}^{r})$ for each $g(w_{i}^{r})$ by Eq. (6):
$\underset{g(w)\in\mathcal{G}}{arg\,min}\;cos\langle g(w_{i}^{r}),g(w)\rangle.$ (6)
@@
History-Aware Aggregation (HAA)
During the aggregation phase, while reducing the weight of highly stale gradients is necessary, it is crucial to identify contributors that potentially possess more valuable data and allocate appropriate additional weights to them. In light of this, $\mathbf{HAA}$ introduces a boosting factor to allocate the clients exhibiting superior performance in historical aggregations with higher weights. We combine the boosting factor with the exponential weighting method derived from (Chen, Sun, and Jin 2019) and control the weights of these two components using a constant $\lambda$, as shown in Eq. (8):
$\hat{p}_{i}^{r}=\sum_{i=1}^{K}\left(\left(\frac{e}{2}\right)^{-\tau(w_{i}^{r})}+\lambda\cdot\overline{\boldsymbol{Util}}{(i,r)}\right),$ (8)
where $\tau(w_{i}^{r})$ and $\overline{\boldsymbol{Util}}{(i,r)}$ represent the staleness and the utility value of the $i$-th client in round $r$, respectively. $\tau(w_{i}^{r})\geq 1$ indicates the number of communication rounds experienced by the client from receiving the global model to uploading its update and $\overline{\boldsymbol{Util}}{(i,r)}$ is obtained based on its historical behavior. Then, $\hat{p}_{i}^{r}$ needs to be normalized:
${p}_{i}^{r}=\frac{\hat{p}_{i}^{r}}{\sum_{i=1}^{K}\hat{p}_{i}^{r}}.$ (9)
A pivotal step involves the computation of the utility value $\overline{\boldsymbol{Util}}$, and $\mathbf{HAA}$ obtains it by identifying relatively fresh updates and generating predicted unbiased gradients for client utility evaluation, as outlined below.
Relatively fresh updates.
@@
$1\displaystyle 0\leq\text{index}(g_{L},\mathcal{L})\leq h-$ 1 .
@@
Predicted unbiased gradients.
Since the local gradients in $\mathcal{S}$ are relatively fresh to round $r-h$, their aggregated result has higher credibility to the global gradient in round $r-h$. By performing average aggregation on them, a predicted unbiased gradient $g_{pred}$ can be generated, representing the ideal global gradient obtained in the absence of staleness scenarios in round $r-h$.
@@
Note that the actual global gradient $g_{act}$ in round $r-h$, i.e., $g(w^{w-h})$, is still formed by aggregating those potentially stale local gradients. In fact, it is staleness that leads to the deviation between $g_{act}$ and $g_{pred}$. In order to assess the quality of the local gradients in round $r-h$, we can calculate their similarity with $g_{pred}$ one by one, followed by the corresponding client utility evaluation.
Client utility evaluation.
At the end of each communication round $r$, the server evaluates the historical local gradients (referred to as $g_{his}^{i}$ hereafter for simplicity) participating in the aggregation phase in round $r-h$. $\mathbf{HAA}$ evaluates the utility of the corresponding client using Eqs. (12) - (13).
$\boldsymbol{Util}(i,r)=(cos\langle g_{his}^{i},g_{pred}\rangle-sim_{thr})\cdot p_{his}^{i}\cdot\lvert\mathcal{S}\lvert,$ (12)
where $sim_{thr}$ is a similarity threshold, and $p_{his}^{i}$ is determined based on the actual (not relative) staleness of $g_{his}^{i}$:
@@
According to Eqs. (12) - (13), the sign of $\boldsymbol{Util}(i,r)$ depends on whether the similarity exceeds the threshold (i.e., reward or penalty). If a gradient has a higher staleness during the actual aggregation process, it is expected to exhibit lower similarity with $g_{pred}$, thereby deserving relatively higher rewards or lower penalties, and vice versa. Moreover, the number of gradients in $\mathcal{S}$ reflects the predictive accuracy of $g_{pred}$, thus $\boldsymbol{Util}$ is positively correlated with $\lvert\mathcal{S}|$. To mitigate the influence of data noise, we employ an exponential moving average with a constant $\gamma$ to obtain the average utility $\overline{\boldsymbol{Util}}{(i,r)}$.
$\overline{\boldsymbol{Util}}{(i,r)}=(1-\gamma)\cdot\overline{\boldsymbol{Util}}{(i,r-1)}+\gamma\cdot\boldsymbol{Util}(i,r).$ (14)
By applying Eqs. (9) - (14) to Eq. (8), $\mathbf{HAA}$ employs a multi-dimensional weighting method by leveraging the participants’ historical updates, emphasizing the gradients from clients with more valuable data, even if they are stale.
Algorithm 1 Federated Historical Learning ((((FedHist))))
1:communication rounds $R$, number of gradients received by server $K$, learning rate $\eta$, length of the server-side buffer $h$, constant $\alpha,\lambda,\gamma,\mu$.
2:Optimal solution $w^{*}$.
3:Initialization.
4:for each communication round $r$ = $1,\ldots,R$ do
5:     for $i=1$ to $K$ do in parallel
6:         Receive $\{g(w_{i}^{r})$$,\tau(w_{i}^{r}\}$ from the $i$-th client.
7:         $\mathcal{L}[h-1]\leftarrow\mathcal{L}[h-1]+[g(w_{i}^{r})]$
8:         if $r\textgreater h$ then $\triangleright$ EGS
9:              Calculate $g(w_{i,co}^{r})$ according to Eq. (6).
10:              $\tilde{g}(w_{i}^{r})\leftarrow g(w_{i}^{r})+\alpha\cdot g(w_{i,co}^{r})$.
11:         else
12:              $\tilde{g}(w_{i}^{r})\leftarrow g(w_{i}^{r})$.
13:         end if
14:     end for
15:     Calculate $p_{i}^{r}$ for $i=1$ to $K$ by Eqs.(8) - (9)$\triangleright$ HAA
16:     $g(w^{r})\leftarrow\sum_{i=1}^{K}p_{i}^{r}g(w_{i}^{r})$.
17:     Adjust the $\ell_{2}$-norm of ${g}(w^{r})$ by Eq. (15)$\triangleright$ INA
18:     Update global model $w^{r}$ by Eq. (4).
19:     Distribute $w^{r}$ to the participants.
20:     Calculate $g_{pred}$ by Eqs. (10) - (11).
21:     Update $\overline{\boldsymbol{Util}}$ by Eqs. (12) - (14).
22:     $\mathcal{L}\leftarrow\mathcal{L}[1:]+[\,[\;]\,]$$\triangleright$ Update the buffer
23:     $\mathcal{G}\leftarrow\mathcal{G}[1:]+[{g}(w^{r})]$$\triangleright$ Cache the global model
24:end for
Intelligent $\ell_{2}$-Norm Amplification (INA)
Due to the impact of Non-IID data, notable biases exist among the local gradients, thus the aggregated gradient typically exhibits a relatively lower $\ell_{2}$-norm, leading to quite limited improvements at each iteration and significantly hindering the convergence speed. An intuitive approach is to make the $\ell_{2}$-norm of the aggregated gradient comparable to that of the local gradients involved in the aggregation. Thus we enhance the convergence speed and stability by Eq. (15):
$\|g(w^{r})\|_{2}=\frac{1-\mu\cdot r}{K}\sum_{i=1}^{K}\|g(w_{i}^{r})\|_{2},$ (15)

### v2

# Knowledge Rumination for Client Utility Evaluation in Heterogeneous Federated Learning

@@ Abstract
Federated Learning (FL) allows several clients to cooperatively train machine learning models without disclosing the raw data. In practical applications, asynchronous FL (AFL) can address the straggler effect compared to synchronous FL. However, Non-IID data and stale models pose significant challenges to AFL, as they can diminish the practicality of the global model and even lead to training failures. In this work, we propose a novel AFL framework called Federated Historical Learning (FedHist), which effectively addresses the challenges posed by both Non-IID data and gradient staleness based on the concept of knowledge rumination. FedHist enhances the stability of local gradients by performing weighted fusion with historical global gradients cached on the server. Relying on hindsight, it assigns aggregation weights to each participant in a multi-dimensional manner during each communication round. To further enhance the efficiency and stability of the training process, we introduce an intelligent $\ell_{2}$-norm amplification scheme, which dynamically regulates the learning progress based on the $\ell_{2}$-norms of the submitted gradients. Extensive experiments indicate FedHist outperforms state-of-the-art methods in terms of convergence performance and test accuracy.
Index Terms:
federated learning, stale gradients, non-iid data
I Introduction
As a distributed machine learning paradigm, Federated learning (FL) [1] has garnered widespread attention in recent years due to its ability to protect the privacy of end-user data. In a classical FL workflow, each client uses its private data to train a local model and then uploads the gradient to a central server. The central server aggregates the collected local gradients into a global gradient to update the global model, which is then sent back to the clients for subsequent training rounds. The entire process does not leak the raw data. As such, FL effectively safeguards data privacy and provides unique advantages in various domains including healthcare [2], autonomous driving [3], and mobile keyboard prediction [4]. A well-known FL aggregation algorithm is federated averaging (FedAvg) [1], where the server aggregates updates from a selected group of local clients and computes the global model by applying a weighted averaging strategy. Due to the system and statistical heterogeneity among devices, this approach often encounters the straggler effect [5], which severely hampers the training progress. In response, [6] introduced asynchronous FL (AFL), where the server no longer waits for all clients but can promptly update the global model upon receiving a single local gradient. However, AFL may lead to global model bias towards fast clients, and the frequent data transmissions may result in server crashes [7]. Some studies have tried to seek a compromise between synchronous and asynchronous approaches [8, 9], where K-async FL [9] has demonstrated promising results, particularly in highly heterogeneous environments. In K-async FL, the server initiates aggregation as soon as it receives the first $K$ local gradients, while the remaining clients continue training.
@@ I Introduction
Despite the favorable performance of K-async FL, it is still constrained by two challenges, as shown in Figure 1. On the one hand, the presence of non-independent and identically distributed (Non-IID) data poses a significant challenge in training a well-performing global model [10]. On the other hand, the existence of stale local gradients often affects the utility of the global gradient, even leading to divergence [11]. Currently, it is common practice to adopt strategies such as variance reduction [12, 13], regularization [14], and momentum [15, 16] to accommodate heterogeneous scenarios. As for the latter, the mainstream approach involves assigning different weights to local gradients based on their staleness [17, 18] to alleviate the weight divergence. While some algorithms can address both aforementioned issues simultaneously, they either rely on a sole weighted criterion [9] leading to a biased allocation of weights, or necessitate specific prior knowledge that is not readily accessible such as the grouping of clients based on varying computational capacities [19], leading to challenges in practical application.
@@
To address these two challenges, drawing inspiration from the concept of Knowledge Rumination, we propose FedHist (the framework is illustrated in Figure 2), a novel K-async FL framework that is capable of handling Non-IID data and stale gradients simultaneously. Knowledge Rumination refers to a behavior, akin to certain animals regurgitating food from the stomach back to the oral cavity for rechewing, aimed at fully absorbing nutrients. Specifically, we leverage a historical gradient buffer at the server-side to facilitate knowledge sharing across clients, enhancing the adaptability to Non-IID data. And we use the historical updates of clients to evaluate their utilities, as a factor for determining their weights during the near future aggregation stage. It is noteworthy that the primary distinction of FedHist from previous works lies in our innovative use of historical updates to evaluate client utility, resolving asynchronous issues from a synchronous perspective. Additionally, we investigate factors that influence convergence and propose a strategy for amplifying the $\ell_{2}$-norms of aggregated gradients adaptively to accelerate the convergence of K-async FL. More importantly, FedHist does not require any prior knowledge about the clients, such as computational capabilities, and maintains the same level of privacy as FedAvg. In summary, FedHist effectively addresses the challenges posed by Non-IID data and stale gradients in AFL. Our main contributions can be summarized as follows:
@@
  • We introduce a gradient buffer to achieve history-aware aggregation on the basis of improving the stability of local gradients and propose a $\ell_{2}$-norm amplification strategy. To the best of our knowledge, we are the first to introduce the concept of knowledge rumination for client utility evaluation, thereby enhancing the effectiveness.
  • Extensive experiments on public datasets demonstrate that FedHist outperforms state-of-the-art methods and effectively promotes fairness in the training process.
Figure 2: The overall framework of FedHist. It comprises three components: EGS, HAA, and INA.
II Methodology
II-A Problem Statement
Suppose there are $N$ clients, denoted as $c_{1},c_{2},\ldots,c_{N}$, and each client $c_{i}$ trains a model using its local dataset $\mathcal{D}_{i}=\{\xi_{i,1},\xi_{i,2},\cdots,\xi_{i,\lvert{\mathcal{D}_{i}}\rvert}\}$, where $\xi_{i,j}$ represents the $j$-th data point $\{x_{j},y_{j}\}$ on client $i$. Our objective is to collaboratively train a global model $w$ using the entire dataset $\mathcal{D}=\bigcup_{i\in N}\mathcal{D}_{i}$ without compromising the privacy of the raw data. For client $c_{i}$, its empirical loss is defined as:
@@ II-A Problem Statement
where $w$ is a parameter vector, which is typically synchronized with the server at the beginning of each round, and $f(\cdot)$ is a loss function. Formally, the objective is to solve:
@@
where $K$ represents the number of participating clients in each communication round. In round $r+1$, client $c_{i}$ updates the parameter vector $w_{i}^{r}$ through mini-batch stochastic gradient descent (SGD) to minimize the objective function:
@@
where $w_{i}^{r+1}$ represents the model of client $c_{i}$ in round $r+1$, $w^{r}$ denotes the global model in round $r$, and $\eta_{i}$ is the learning rate of client $c_{i}$. We use $g(w_{i}^{r})$ to denote the gradient submitted by $i$-th client in round $r$. Upon receiving $K$ gradients, the server updates the global model parameters accordingly:
@@
Here, $\eta$ is the learning rate of the server, $p_{i}^{r}$ denotes the weight of $i$-th client in round $r$, and $\sum_{i=1}^{K}p_{i}^{r}=1$.
II-B Enhancement of Gradient Stability (EGS)
FedHist incorporates a buffer at the server-side, comprising $\mathcal{G}$ and $\mathcal{L}$, which respectively cache global gradients and local gradients of past $h$ communication rounds from clients.
The goal of $\mathbf{EGS}$ is to select an appropriate collaboration gradient from $\mathcal{G}$ for each local gradient uploaded from a client, and then execute a weighted aggregation process. Quantitatively measuring the degree of similarity between two gradient vectors $g(w_{i})$ and $g(w_{j})$ can be achieved by computing their cosine similarity, as shown in Eq. (5):
$\mathbf{sim}\langle g(w_{i}),g(w_{j})\rangle=\frac{{g(w_{i})}\cdot{g(w_{j})}}{\|{g(w_{i})}\|\,\|{g(w_{j})}\|}.$ (5)
We select the corresponding collaborative gradient for each local gradient based on the principle of minimizing similarity, as it implies a higher potential of sharing knowledge between two gradients and avoids overfitting [20]. This fine-grained cross-aggregation method mitigates gradient divergence and rectifies bias in local gradients. Thus the objective is to find the optimal collaborative gradient $g(w_{i,co}^{r})$ in $\mathcal{G}$ for each $g(w_{i}^{r})$:
$\underset{g(w)\in\mathcal{G}}{arg\,min}\;\mathbf{sim}\langle g(w_{i}^{r}),g(w)\rangle.$ (6)
@@ II-B Enhancement of Gradient Stability (EGS)
II-C History-Aware Aggregation (HAA)
To promote fairness in FL, a key aspect is to identify potential contributors with more valuable data and allocate appropriate additional weights to them. With this goal in mind, $\mathbf{HAA}$ introduces a boosting factor to allocate higher weights to clients that exhibit outstanding performance in historically weighted aggregations. We combine the boosting factor with the exponential weighting method derived from [17], using $\lambda$ to control the weights of these two components:
$\hat{p}_{i}^{r}=\sum_{i=1}^{K}\left(\left(\frac{e}{2}\right)^{-\tau(w_{i}^{r})}+\lambda\cdot\overline{\bm{Util}}{(i,r)}\right),$ (8)
where $\tau(w_{i}^{r})$ and $\overline{\bm{Util}}{(i,r)}$ represent the staleness and utility value of the $i$-th client in round $r$, respectively. $\tau(w_{i}^{r})\geq 1$ indicates the number of communication rounds have passed from the time the client $i$ received the last global model to the moment it starts uploading its new update and $\overline{\bm{Util}}{(i,r)}$ is obtained based on its historical behavior. And $\hat{p}_{i}^{r}$ needs to be normalized: ${p}_{i}^{r}={\hat{p}_{i}^{r}}/{\sum_{i=1}^{K}\hat{p}_{i}^{r}}.$
On the whole, building upon the concept of knowledge rumination, $\mathbf{HAA}$ employs the reminiscence of past data to generate unbiased gradients for prediction. It then compares the predicted values with the ground truth to calculate the client utility $\overline{\bm{Util}}$. In fact, when $N$ is much higher than $K$, the lower bound approximation for the average staleness is close to $\frac{N}{2K}$ [9], which demonstrates that the local gradients uploaded by clients undergo a fixed number of rounds $\left(N/(2K)\right)$ of historical backtracking, resulting in obtaining corresponding relatively fresh updates.
Relatively fresh updates
@@ II-C History-Aware Aggregation (HAA)
$1\leq\text{index}(g_{L},\mathcal{L})\leq h.$
@@
Predicted unbiased gradients
Since local gradients in $\mathcal{S}$ are relatively fresh to round $r-h$, their aggregated result has higher credibility to the global gradient in round $r-h$. By performing aggregation on them, a predicted unbiased gradient $g_{pred}$ can be generated, representing the ideal global gradient obtained in the absence of staleness in round $r-h$.
@@
Note that the actual global gradient $g_{act}$ in round $r-h$, i.e., $g(w^{w-h})$, is still formed by aggregating those potentially stale local gradients. In fact, it is staleness that leads to the deviation between $g_{act}$ and $g_{pred}$. In order to assess the quality of the local gradients in round $r-h$, we can replay the previous knowledge to calculate their similarity with $g_{pred}$, followed by the corresponding client utility evaluation.
Client utility evaluation
At the end of each communication round, the server evaluates the historical local gradients (referred to as $g_{his}^{i}$ hereafter for simplicity) participating in the aggregation phase in round $r-h$. $\mathbf{HAA}$ evaluates the utility of the corresponding clients using Eqs. (11) - (12).
$\bm{Util}(i,r)=(\mathbf{sim}\langle g_{his}^{i},g_{pred}\rangle-{sim}_{T})\cdot p_{his}^{i}\cdot\lvert\mathcal{S}\lvert,$ (11)
where ${sim}_{T}$ is a similarity threshold, and $p_{his}^{i}$ is determined based on the actual (not relative) staleness of $g_{his}^{i}$:
@@
According to Eqs. (11) - (12), the sign of $\bm{Util}(i,r)$ depends on whether the similarity exceeds the threshold (i.e., reward or penalty). If a gradient has a higher staleness during the actual aggregation process, it is expected to exhibit lower similarity with $g_{pred}$, thereby deserving relatively higher rewards or lower penalties, and vice versa. Moreover, the number of gradients in $\mathcal{S}$ reflects the predictive accuracy of $g_{pred}$, thus $\bm{Util}$ is positively correlated with $\lvert\mathcal{S}|$. To mitigate the influence of data noise, we employ an exponential moving average with a constant $\gamma$ to obtain the average utility $\overline{\bm{Util}}{(i,r)}$.
$\overline{\bm{Util}}{(i,r)}=(1-\gamma)\cdot\overline{\bm{Util}}{(i,r-1)}+\gamma\cdot\bm{Util}(i,r).$ (13)
By applying Eqs. (9) - (13) to Eq. (8), $\mathbf{HAA}$ employs a multidimensional weighting approach based on the replay of historical gradients to explore stale yet more valuable gradient information, thereby promoting fairness.
II-D Intelligent $\ell_{2}$-Norm Amplification (INA)
Key observation. In K-async FL, the aggregation of local gradients often leads to a significant reduction in the $\ell_{2}$-norm of the aggregated result, which severely hinders the convergence process. Therefore, the directionality of the aggregated result holds more reference value, while its magnitude does not. Unfortunately, this observation is unexpectedly overlooked in many related methods [17]. To address this, we re-adjust the global gradients:
$g(w^{r})=\frac{\sum_{i=1}^{K}g(w_{i}^{r})}{\|\sum_{i=1}^{K}g(w_{i}^{r})\|_{2}}\cdot\frac{1}{K}\sum_{i=1}^{K}\|g(w_{i}^{r})\|_{2}.$ (14)